# Retrieval Enhanced Zero-Shot Video Captioning

### v1

@@ Abstract.
Despite the significant progress of fully-supervised video captioning, zero-shot methods remain much less explored. In this paper, we propose to take advantage of existing pre-trained large-scale vision and language models to directly generate captions with test-time adaptation. Specifically, we bridge video and text using three key models: a general video understanding model XCLIP, a general image understanding model CLIP, and a text generation model GPT-2, due to their source-code availability. The main challenge is how to enable the text generation model to be sufficiently aware of the content in a given video so as to generate corresponding captions.To address this problem, we propose using learnable tokens as a communication medium between frozen GPT-2 and frozen XCLIP as well as frozen CLIP. Differing from the conventional way to train these tokens with training data, we update these tokens with pseudo-targets of the inference data under several carefully crafted loss functions which enable the tokens to absorb video information catered for GPT-2. This procedure can be done in just a few iterations (we use 16 iterations in the experiments) and does not require ground truth data. Extensive experimental results on three widely used datasets, MSR-VTT, MSVD, and VATEX, show 4$\sim$20improvements in terms of the main metric CIDEr compared to the existing state-of-the-art methods.
*Laiyun Qing is the corresponding author
copyright: none
1. Introduction
@@
Video captioning, aiming to automatically describe the video content in one sentence, is one of the traditional vision-language tasks. It plays a crucial role in numerous applications, such as video title generation (Zhang et al., 2020), blind assistance (Huo et al., 2021), and video search (Nie et al., 2022). Although current video captioning research makes remarkable progress on full-supervision algorithms, exploration in the area of zero-shot learning still remains limited.
@@
Recently, with the development of pre-trained foundation models in vision and language areas, their combination has shown outstanding zero-shot learning capabilities. For example, for the text-based question-answering task, pre-trained retrieval models have been integrated with pre-trained large language models (PLMs) to improve the quality of generated answers. An example is shown in Fig. 1 (a), where the retriever takes the question “The 2024 Olympics will be held” as a query to retrieve relevant knowledge “The 2024 Summer Olympics, known as Paris 2024, will be hosted in Paris, France, from 26 July to 11 August.”. By concatenating it with the question as a prompt for the model, we get the output “Paris, France, from 26 July to 11 August 2024”. However, a simple application of this paradigm does not work well for multi-modal visual description tasks (e.g., video captioning). For example, as shown in Fig. 1 (b), the video query retrieves the related sentence “Cutting a piece of bread”; then together with the hard prompt “Video showing” serving as the input of PLM, the PLM outputs “how to cut”, which does not meet the expectation of video captioning. The main reasons lie in two aspects: (1) Weak semantic connection: The semantic connection between the retrieved text information and the prompt “video showing” is not as strong as that between questions and answers in QA tasks. This hampers PLMs to generate expected outputs. (2) Improper prompts: Prompt engineering itself is a challenging task, letting alone to additionally take diverse retrieval results into consideration.
To address the above-mentioned problems, we propose to leverage learnable tokens as a connection medium among different foundation models. These tokens should indicate to PLMs what type of output is expected and convey video information to PLMs. Moreover, we expect the combination of these models is able to quickly adapt to the task, preferably without the need of training. To meet the quick adaptation requirement, we propose to update these tokens during inference but with pseudo-targets, which gets rid of the conventional training of soft prompts. Our experiments show that the whole framework can well adapt to the video captioning task after just a few of iterations (we use 16 iterations). To meet the former output type and content awareness requirements, we design several loss functions aligning the generated text and pseudo-target as well as retrieved sentences.
@@
Figure 1. Traditional retrieval-augmented-generation paradigm performs well on pure natural language tasks, such as text-based question answering. However, its performance sucks when directly apply to multi-modality tasks, such as video captioning.
An example is shown in Fig. 1(c), where our method leads to better video captions.
@@
  • To the best of our knowledge, we propose the first fast adaptation framework that combines a pre-trained video-text retrieval model, an image-text matching model, and a large language generation model for video captioning.
@@
  • The extensive experimental results on three widely-used datasets show that our method outperforms several state-of-the-art zero-shot video captioning approaches with 4$\sim$20improvement in terms of the main metric CIDEr.
Figure 2. The workflow of the proposed framework, which consists of frozen pre-trained foundation vision and language models as well as trainable tokens. Unlike conventional soft prompt training, we update these tokens directly during inference with pseudo-target. In this way, the frozen foundation models can quickly adapt to the video captioning task for zero-shot application. $\mathcal{L}_{language}$ is computed for word probability distributions when using (i.e. writing) and not using (i.e. reading) soft prompt. $\mathcal{L}_{vision}$ calculates a matching score between the current generated text and visual information leveraging CLIP based on keyframes. $\mathcal{L}_{retrieval}^{S}$ measures a matching score between the current generated text and retrieved sentences using AnglE, and $\mathcal{L}_{retrieval}^{W}$ focuses on the high-frequency words of them. The above optimizations take place at each step of the autoregressive process.
2. Related work
@@
Video Captioning Recently, a series of full-supervision video captioning research have made significant progress. However, fewer efforts have been made on the video captioning task under the zero-shot setting (Tewel et al., 2022b, a; Su et al., 2022; Jo et al., 2023). ZeroCap (Tewel et al., 2022b) uses a visual-language model to guide a language model to describe the given image via the designed CLIP-loss. EPT (Tewel et al., 2022a) further extends it to video captioning using the proposed CLIP-loss to optimize the pseudo-tokens. To improve inference speed, MAGIC (Su et al., 2022) optimizes the decoding strategy of language models and employs CLIP scores to adjust PLM logits for better alignment with image correspondence. DeCap (Li et al., 2023) and MultiCapCLIP (Yang et al., 2023) also leverage the visual-language alignment ability of CLIP for zero-shot video captioning. They first utilize text-only data to train a decoder from scratch and then input video to generate captions at the inference stage. Although the above-mentioned methods improve zero-shot caption quality, they still face the challenge of understanding video content. To tackle this problem, we propose a retrieval-augmented video captioning framework that retrieves video-related knowledge to deepen the understanding of the video.
@@
Video-Text Retrieval Video-text retrieval (VTR) aims to find the most relevant text based on a video query. To retrieve the correct video candidates given a text query, recent works (Luo et al., 2022; Gorti et al., 2022; Liu et al., 2022; Zou et al., 2022; Ma et al., 2022; Ni et al., 2022) focus on the cross-modal alignment between the related videos and text samples. Most of them can be classified into coarse-grained alignment or fine-grained alignment. The coarse-grained alignment methods mainly leverage the frame-level or video-level visual features to perform alignment between the video and query features. Clip4clip (Luo et al., 2022) transfers the image-text knowledge of the CLIP model to the VTR task in a coarse manner. X-Pool (Gorti et al., 2022) designs a cross-modal attention model that performs coarse-grained alignment between a text and the frames of a video. TS2-Net (Liu et al., 2022) introduces the token shift and selection transformer to better model video representation in both temporal and spatial dimensions. However, the coarse-grained alignment is not suitable for capturing detailed correspondence between texts and videos. To overcome this challenge, TokenFlow (Zou et al., 2022) presents a fine-grained alignment for token-wise similarity calculation. X-CLIP (Ma et al., 2022) adopts multi-grained contrastive learning to boost video-text retrieval performance, covering fine-grained, coarse-grained, and cross-grained contrast. Another X-CLIP (Ni et al., 2022) extends CLIP for general video-language understanding using the Kinetics-600 (Li et al., 2020) dataset. In this work, we adopt (Ni et al., 2022) as our retriever as its training data has no overlap with video captioning datasets.
@@
Retrieval-Augmented Generation Retrieval-Augmented Generation (RAG) augments the generation ability of language models with retrieved knowledge. Generally, the knowledge retrieved is used as a prefix input to the pre-trained language model (PLM), prompting PLM to generate more accurate answers. REALM (Guu et al., 2020) employs a latent knowledge retriever to boost language model pre-training, which allows capturing more real-world knowledge in an interpretable manner. ORQA (Lee et al., 2019) pre-trains the retriever with an unsupervised inverse cloze task and learns to retrieve evidence from an open corpus. DPR (Karpukhin et al., 2020) implements a dense embedding retriever to better select candidate contexts using only pairs of questions and answers, without additional pretraining. RCG (Zhang et al., 2021) plugs a cross-modal retriever to provide useful hints and employs a dynamic copy-mechanism generator for better generation. Recently, with the development of large language models (LLMs), a series of RAG-related LLM approaches are proposed. RECOMP (Xu et al., 2023) uses a summarization model to compress retrieved documents into a textual summary before prompting the LLM to generate the output. Chen et al. establish the first benchmark to evaluate robustness-related capabilities for retrieval augmented generation of LLMs. Asai et al. further introduce a new framework SELF-RAG, which trains an arbitrary LM to generate text with reflection tokens by integrating them as the next token prediction from the expanded model vocabulary.
@@
3. Method
@@
As shown in Fig. 2, our framework contains several pre-trained frozen models: XCLIP (Ni et al., 2022) and CLIP (Radford et al., 2021) provide video-level and frame-level video understanding; GPT-2 (Radford et al., 2019) provide the text generation capability; AnglE (Li and Li, 2023) is used to calculate text-text similarity. In a big view, according to their functions for the video captioning task, these pre-trained models can be grouped into a video-text retrieval module and an augmented caption generation module. Specifically, XCLIP serves as the video-text retrieval module, providing a basic understanding of the video. It retrieves video content-related sentences, and we highlight its high-frequency words, as detailed in Section 3.1. The rest pre-trained models compose the augment caption generation module, which utilizes several learnable tokens to connect all these pre-trained models, absorb video information, and serve as proper prompts for text generation as detailed in Section 3.2. In Section 3.3, we provide more details of the inference process.
3.1. Video-text Retrieval
@@
The video-text retrieval module adopts a multi-modal retriever XCLIP to retrieve video-related descriptions from a text corpus. XCLIP is pre-trained on the Kinetics-600 (Li et al., 2020) dataset and has the zero-shot ability for video classification and video-text retrieval tasks. Traditional retrievers consistently fetch the top-$k$ ranked sentences. However, this retrieval strategy faces two challenges for video captioning. First, choosing the hyper-parameter $k$ is challenging since the optimal $k$ varies across different videos. Second, retrieving a fixed number of $k$ sentences may introduce noise, as not all of these sentences are closely related to the video.
@@
To mitigate the aforementioned issues, we additionally sample top-$L$ high-frequency words (limited to nouns and verbs) from the retrieved sentences. In the next section, we detail how retrieved sentences and sampled high-frequency words are utilized to guide caption generation.
3.2. Augmented Caption Generation
@@
With the help of the video-text retrieval module, the augmented caption generation module uses the retrieved sentences and high-frequency words to improve the quality of the generated captions. This module contains three main components: the large language model GPT-2, the image-text matching model CLIP, and the sentence-alignment model. GPT-2 is trained on the large dataset WebText collected from a wide range of text data online. It is composed of several Transformer layers, which can cache the previous key and value tokens to infer the next token autoregressively. Therefore, GPT-2 is capable of generating text based on preceding content. The CLIP model is pre-trained on 400 million image-text pairs, which plays the vision-language alignment role to map the image and sentence into a unified semantic space. AnglE is a state-of-the-art model for calculating semantic textual similarity, which is able to be aware of differences among texts.
@@
As shown in Fig. 2, all model parameters are fixed, and only the learnable soft prompts are optimized during the inference process to guide GPT-2 to generate text in line with the content of the given video. Take the $i_{th}$ step as an example, the text generation can be described as:
(1) $w_{i+1}=PLM(w_{soft},w_{hard},s_{i})$
@@
where $w_{i+1}$ denotes the next to be generated word and $s_{i}$ denotes all the existing words generated in previous steps. $w_{soft}$ is the set of learnable tokens that can be updated by the guidance during the generation process. $w_{hard}$ is the hard prompt, which is randomly sampled from a predefined prompt set such as “Video showing”, “Video describes”. To utilize the retrieved sentence, we propose retrieval-based loss $\mathcal{L}_{retrieval}$ to transfer the retrieved video content into the the above auto-regression process. It consists of two granularities: the sentence-granularity and the word-granularity. Specifically, the $\mathcal{L}_{retrieval}$ is formulated as
(2) ${}\mathcal{L}_{retrieval}^{S}=-\sum_{\mathrm{k}=1}^{\mathrm{N}}\mathrm{p}\left(\mathrm{w}_{k}\right)\log\left(\mathrm{q}\left(\mathrm{w}_{k}\right)\right)$
@@
where $\mathrm{q}\left(\mathrm{w}_{k}\right)$ is the probability of the $k$-th word in the vocabulary, and $\mathrm{p}(\cdot)$ denotes the pseudo-target probability. Specifically, $\mathrm{p}\left(\mathrm{w}_{k}\right)$ can be computed as
(3) $\mathrm{p}\left(\mathrm{w}_{k}\right)=\frac{1}{K}\sum_{R\in\mathcal{R}}\operatorname{AnglE}\left(R,s_{k}\right)$
@@
where $K$ denotes the number of retrieved sentences, $\mathrm{p}\left(\mathrm{w}_{k}\right)$ is the average of $\operatorname{AnglE}$’s alignment scores between $s_{k}$ and all the retrieved sentence in $\mathcal{R}$. This loss adjusts learnable tokens so as to prompt GPT-2 to generate texts better aligned with the retrieved description of the given video. For efficiency, we only calculate the scores for the top 100 candidate words based on the original probability distribution obtained by GPT-2. On the other hand, considering that the retrieved sentences might contain noise, we propose a word-granularity loss, which focuses on the high-frequency information in the retrieved sentences. The word-granularity retrieval loss is formulated as
(4) ${}\mathcal{L}_{retrieval}^{W}=-\sum_{\mathrm{k}=1}^{\mathrm{N}}\mathrm{\hat{p}}\left(\mathrm{w}_{k}\right)\log\left(\mathrm{q}\left(\mathrm{w}_{k}\right)\right)$
(5) $\mathrm{p}\left(\mathrm{w}_{k}\right)=\frac{1}{L}\sum_{W\in\mathcal{W}}\operatorname{AnglE}\left(W,s_{k}\right)$
@@
where $\mathcal{W}$ denotes all the high-frequency words in the retrieved sentence. $\mathrm{p}\left(\mathrm{w}_{k}\right)$ is the average of $\operatorname{AnglE}$’s alignment scores between $s_{k}$ and all the sampled high-frequency words in $\mathcal{W}$.
@@
Inspired by (Tewel et al., 2022a, b; Jo et al., 2023), we also employ two losses to optimize the augment caption generation module. The first loss aims to reduce the gap between the generated caption and the corresponding video. It can be described as:
@@
(6) $\mathcal{L}_{vision}=-\sum_{\mathrm{k}=1}^{\mathrm{N}}(\frac{1}{T}\sum_{F\in\mathcal{F}}\operatorname{CLIP}\left(F,s_{k}\right))\log\left(\mathrm{q}\left(\mathrm{w}_{k}\right)\right)$
@@
where $T$ is the total number of sampled video frames and we use these video frames to represent a video. CLIP calculates the average matching score between each sampled frame $F$ in $\mathcal{F}$ and $s_{k}$.
@@
(7) $\mathcal{L}_{\text{language}}=\text{CE}\left(\mathrm{q}(w_{i+1}),\mathrm{q}(\hat{w}_{i+1})\right)$
(8) $\hat{w}_{i+1}=PLM(w_{hard},s_{i})$
@@
3.3. Inference
In our zero-shot setting, we directly perform testing on the test set, without any videos from the training set for training. The learnable tokens are optimized by the total loss:
(9) $\mathcal{L}_{total}=\lambda_{l}\mathcal{L}_{language}+\lambda_{s}\mathcal{L}_{retrieval}^{S}+\lambda_{w}\mathcal{L}_{retrieval}^{W}+\lambda_{v}\mathcal{L}_{vision}$

### v2

# RETTA: Retrieval-Enhanced Test-Time Adaptation for Zero-Shot Video Captioning

@@ Abstract
Despite the significant progress of fully-supervised video captioning, zero-shot methods remain much less explored. In this paper, we propose a novel zero-shot video captioning framework named Retrieval-Enhanced Test-Time Adaptation (RETTA), which takes advantage of existing pre-trained large-scale vision and language models to directly generate captions with test-time adaptation. Specifically, we bridge video and text using four key models: a general video-text retrieval model XCLIP, a general image-text matching model CLIP, a text alignment model AnglE, and a text generation model GPT-2, due to their source-code availability. The main challenge is how to enable the text generation model to be sufficiently aware of the content in a given video so as to generate corresponding captions.To address this problem, we propose using learnable tokens as a communication medium among these four frozen models GPT-2, XCLIP, CLIP, and AnglE. Different from the conventional way that trains these tokens with training data, we propose to learn these tokens with soft targets of the inference data under several carefully crafted loss functions, which enable the tokens to absorb video information catered for GPT-2. This procedure can be efficiently done in just a few iterations (we use 16 iterations in the experiments) and does not require ground truth data. Extensive experimental results on three widely used datasets, MSR-VTT, MSVD, and VATEX, show absolute 5.1$\sim$32.4improvements in terms of the main metric CIDEr compared to several state-of-the-art zero-shot video captioning methods.
keywords:
[…]
1 Introduction
@@
Video captioning, aiming to automatically describe the video content in one sentence, is one of the traditional vision-language tasks. It plays a crucial role in numerous applications, such as video title generation [1], blind assistance [2], and video search [3]. Although current video captioning research makes remarkable progress on full-supervision algorithms, exploration in the area of zero-shot learning still remains limited.
@@
Recently, with the development of pre-trained foundation models in vision and language areas, their combination has shown outstanding zero-shot learning capabilities. For example, for the text-based question-answering task, pre-trained retrieval models have been integrated with Pre-trained Large language Models (PLMs) to improve the quality of generated answers. An example is shown in Fig. 1 (a), where the retriever takes the question “The 2024 Olympics will be held” as a query to retrieve relevant knowledge “The 2024 Summer Olympics, known as Paris 2024, will be hosted in Paris, France, from 26 July to 11 August.”. By concatenating it with the question as a prompt for the model, we get the output “Paris, France, from 26 July to 11 August 2024”.
@@
Figure 1: Traditional retrieval-augmented-generation paradigm performs well on pure natural language tasks, such as text-based question answering. However, its performance sucks when directly apply to multi-modality tasks, such as video captioning.
However, a simple application of this paradigm does not work well for multi-modal visual description tasks (e.g., video captioning). For example, as shown in Fig. 1 (b), the video query retrieves the related sentence “Cutting a piece of bread”; then together with the hard prompt “Video showing” serving as the input of PLM, the PLM outputs “how to cut”, which does not meet the expectation of video captioning. The main reasons lie in two aspects: (1) Weak semantic connection: The semantic connection between the prompt “video showing” and the retrieved text is not as strong as that between questions and answers in QA tasks. This hampers PLMs to generate expected outputs. (2) Improper prompts: Prompt engineering itself is a challenging task, letting alone to additionally take diverse retrieved text into consideration.
To address the above-mentioned problems, we propose to leverage learnable tokens as a connection medium among different foundation models. These tokens should convey video information to PLMs and indicate to PLMs what type of output is expected. Moreover, we expect the combination of these pre-trained models is able to quickly adapt to the task, preferably without the need of training. To meet the quick adaptation requirement, we propose to update these tokens during inference but with soft targets, which gets rid of the conventional training of soft prompts. Our experiments show that the whole framework can well adapt to the video captioning task after just a few of iterations (we use 16 iterations). To meet the former content awareness and output type requirements, we design several loss functions aligning the generated text and soft targets. Specifically, the soft targets are calculated by retrieved text and candidate sentences to guide the choice of candidate words at each time step. This is because soft targets can help the model consider multiple potential tokens in text generation, boosting the diversity of the generated results and improving generalization. An example is shown in Fig. 1(c), where our method leads to better video captions.
@@
  • To the best of our knowledge, we propose the first fast adaptation framework that combines a pre-trained video-text retrieval model, an image-text matching model, a text alignment model, and a language generation model for video captioning.
@@
  • Extensive experimental results on three widely-used datasets show that our method outperforms several state-of-the-art zero-shot video captioning approaches with absolute 5.1$\sim$32.4improvement in terms of the main metric CIDEr.
2 Related work
@@
Video Captioning Recently, a series of full-supervision video captioning research [4, 5, 6, 7, 8, 9, 10, 11, 12] have made significant progress. However, fewer efforts have been made on the video captioning task under the zero-shot setting [13, 14, 15, 16]. ZeroCap [13] uses a visual-language model to guide a language model to describe the given image via the designed CLIP-loss. EPT [14] further extends it to video captioning using the proposed CLIP-loss to optimize the pseudo-tokens. To improve inference speed, MAGIC [15] optimizes the decoding strategy of language models and employs CLIP scores to adjust PLM logits for better alignment with image correspondence. DeCap [17] and MultiCapCLIP [18] also leverage the visual-language alignment ability of CLIP for zero-shot video captioning. They first utilize text-only data to train a decoder from scratch and then input video to generate captions at the inference stage. Although the above-mentioned methods improve zero-shot caption quality, they still face the challenge of understanding video content. To tackle this problem, we propose a retrieval-augmented video captioning framework that retrieves video-related knowledge to deepen the understanding of the video.
@@
Video-Text Retrieval Video-text retrieval (VTR) aims to find the most relevant text based on a video query. To retrieve the correct video candidates given a text query, recent works [19, 20, 21, 22, 23, 24] focus on the cross-modal alignment between the related videos and text samples. Most of them can be classified into coarse-grained alignment or fine-grained alignment. The coarse-grained alignment methods mainly leverage the frame-level or video-level visual features to perform alignment between the video and query features. Clip4clip [19] transfers the image-text knowledge of the CLIP model to the VTR task in a coarse manner. X-Pool [20] designs a cross-modal attention model that performs coarse-grained alignment between a text and the frames of a video. TS2-Net [21] introduces the token shift and selection transformer to better model video representation in both temporal and spatial dimensions. However, the coarse-grained alignment is not suitable for capturing detailed correspondence between texts and videos. To overcome this challenge, TokenFlow [22] presents a fine-grained alignment for token-wise similarity calculation. X-CLIP [23] adopts multi-grained contrastive learning to boost video-text retrieval performance, covering fine-grained, coarse-grained, and cross-grained contrast. Another X-CLIP [24] extends CLIP for general video-language understanding using the Kinetics-600 [25] dataset. In this work, we adopt [24] as our retriever as its training data has no overlap with video captioning datasets.
@@
Retrieval-Augmented Generation Retrieval-Augmented Generation (RAG) augments the generation ability of language models with retrieved knowledge. Generally, the knowledge retrieved is used as a prefix input to the pre-trained language model (PLM), prompting PLM to generate more accurate answers. REALM [26] employs a latent knowledge retriever to boost language model pre-training, which allows capturing more real-world knowledge in an interpretable manner. ORQA [27] pre-trains the retriever with an unsupervised inverse cloze task and learns to retrieve evidence from an open corpus. DPR [28] implements a dense embedding retriever to better select candidate contexts using only pairs of questions and answers, without additional pretraining. RCG [29] plugs a cross-modal retriever to provide useful hints and employs a dynamic copy-mechanism generator for better generation. Recently, with the development of large language models (LLMs), a series of RAG-related LLM approaches are proposed. RECOMP [30] uses a summarization model to compress retrieved documents into a textual summary before prompting the LLM to generate the output. Chen et al. establish the first benchmark to evaluate robustness-related capabilities for retrieval augmented generation of LLMs. Asai et al. further introduce a new framework SELF-RAG, which trains an arbitrary LM to generate text with reflection tokens by integrating them as the next token prediction from the expanded model vocabulary.
@@
Figure 2: The workflow of the proposed method, which consists of frozen pre-trained foundation vision and language models as well as trainable tokens. Unlike conventional soft prompt training, we update these tokens directly during inference with soft target. In this way, the frozen foundation models can quickly adapt to the video captioning task for zero-shot application. $\mathcal{L}_{language}$ is computed between token probability distributions generated with (i.e. writing) and without (i.e. reading) the trainable soft prompt. $\mathcal{L}_{vision}$ calculates a matching score between the current generated text and visual information leveraging CLIP based on keyframes. $\mathcal{L}_{retrieval}^{S}$ measures a matching score between the current generated text and retrieved sentences using AnglE, and $\mathcal{L}_{retrieval}^{W}$ focuses on the high-frequency words of them. The above optimizations take place at each step of the autoregressive process.
3 Method
@@
As shown in Fig. 2, our framework contains four pre-trained frozen models: XCLIP [24] and CLIP [31] provide video-level and frame-level video understanding; GPT-2 [32] provide the text generation capability; AnglE [33] is used to calculate text-text similarity. According to their functions for the video captioning task, these pre-trained models can be grouped into a video-text retrieval module and an augmented caption generation module. Specifically, XCLIP serves as the video-text retrieval module, providing a basic understanding of the video. It retrieves video content-related sentences from training text corpus, and we highlight its high-frequency words, as detailed in Section 3.1. The remaining pre-trained models compose the augment caption generation module, which utilizes several test-time learnable tokens to connect all these pre-trained models, absorb video information, and serve as proper prompts for text generation as detailed in Section 3.2. In Section 3.3, we provide more details of the inference process.
3.1 Video-Text Retrieval
@@
The video-text retrieval module adopts a multi-modal retriever XCLIP to retrieve video-related descriptions from a text corpus. XCLIP is pre-trained on the Kinetics-600 [25] dataset and has the zero-shot ability for video classification and video-text retrieval tasks. Traditional retrievers consistently fetch the top-$k$ ranked sentences. However, this retrieval strategy faces two challenges for video captioning. First, choosing the hyper-parameter $k$ is challenging since the optimal $k$ varies across different videos. Second, retrieving a fixed number of $k$ sentences may introduce noise, as not all of these sentences are closely related to the video.
@@
To mitigate the aforementioned issues, we propose to additionally sample top-$L$ high-frequency words (limited to nouns and verbs) from the retrieved sentences, which reduces diversity of sentences. We detail how retrieved sentences and sampled high-frequency words are utilized to guide caption generation in the next section.
3.2 Augmented Caption Generation
@@
With the help of the video-text retrieval module, the augmented caption generation module uses the retrieved sentences and high-frequency words to improve the quality of the generated captions. This module contains three main components: the large language model GPT-2, the image-text matching model CLIP, and the text alignment model AnglE. The out-of-box GPT-2 is pre-trained on the large dataset WebText collected from a wide range of text data online. It is composed of several Transformer layers, which can cache the previous key and value tokens to infer the next token autoregressively. Therefore, GPT-2 is capable of generating text based on preceding content. The CLIP model is pre-trained on 400 million image-text pairs, which plays the vision-language alignment role to map the image and sentence into a unified semantic space. AnglE is a state-of-the-art model for calculating semantic textual similarity, which is able to be aware of differences among texts.
@@
As shown in Fig. 2, all model parameters are frozen, and only the learnable soft prompts are optimized during the inference process with soft targets to guide GPT-2 to generate text in line with the content of the given video. Take the $i-{th}$ step as an example, the text generation can be described as:
$w_{i+1}=\operatorname{PLM}(w_{soft},w_{hard},s_{i})$ (1)
@@
where $w_{i+1}$ denotes the next to be generated word and $s_{i}$ denotes all the existing words generated in previous steps. $w_{soft}$ is the set of learnable tokens that can be updated by the guidance during the generation process. $w_{hard}$ is the hard prompt, which is randomly sampled from a predefined prompt set, such as “Video showing”, “Video of”.
To utilize the retrieved sentences, we transfer them into the above auto-regression process with our proposed retrieval-based loss $\mathcal{L}_{retrieval}$. It consists of two granularities: the sentence-granularity $\mathcal{L}_{retrieval}^{S}$ and the word-granularity $\mathcal{L}_{retrieval}^{W}$ Specifically, the $\mathcal{L}_{retrieval}^{S}$ is formulated as
${}\mathcal{L}_{retrieval}^{S}=-\sum_{\mathrm{k}=1}^{\mathrm{N}}\mathrm{p}\left(\mathrm{w}_{k}\right)\log\left(\mathrm{q}\left(\mathrm{w}_{k}\right)\right)$ (2)
@@
where $\mathrm{q}\left(\mathrm{w}_{k}\right)$ is the probability of the $k$-th word in the vocabulary, and $\mathrm{p}(\cdot)$ denotes the probability of being soft target. $\mathrm{p}\left(\mathrm{w}_{k}\right)$ is computed as
$\mathrm{p}\left(\mathrm{w}_{k}\right)=\frac{1}{K}\sum_{R\in\mathcal{R}}\operatorname{AnglE}\left(R,s_{k}\right)$ (3)
@@
where $K$ denotes the number of retrieved sentences, $s_{k}$ represents the combination of the generated words in previous time steps with the $k$-th word, $\mathrm{p}\left(\mathrm{w}_{k}\right)$ is the average of $\operatorname{AnglE}$’s alignment scores between $s_{k}$ and all the retrieved sentence in $\mathcal{R}$. This loss adjusts learnable tokens so as to prompt GPT-2 to generate texts better reflect the content in the retrieved description of the given video. For efficiency, we only calculate the scores for the top 100 candidate words based on the original probability distribution obtained by GPT-2.
On the other hand, considering that the retrieved sentences might contain noise, we propose a word-granularity loss, which focuses on the high-frequency information in the retrieved sentences. The word-granularity retrieval loss is formulated as
${}\mathcal{L}_{retrieval}^{W}=-\sum_{\mathrm{k}=1}^{\mathrm{N}}\mathrm{\hat{p}}\left(\mathrm{w}_{k}\right)\log\left(\mathrm{q}\left(\mathrm{w}_{k}\right)\right)$ (4)
$\mathrm{\hat{p}}\left(\mathrm{w}_{k}\right)=\frac{1}{L}\sum_{W\in\mathcal{W}}\operatorname{AnglE}\left(W,s_{k}\right)$ (5)
@@
where $\mathcal{W}$ denotes all the high-frequency words in the retrieved sentences. $\mathrm{\hat{p}}\left(\mathrm{w}_{k}\right)$ is the average of $\operatorname{AnglE}$’s alignment scores between $s_{k}$ and all the sampled high-frequency words in $\mathcal{W}$.
@@
Inspired by [13, 14, 16], we also employ two reliable losses to optimize the augment caption generation module. The first loss aims to reduce the gap between the generated caption and the corresponding video:
@@
$\mathcal{L}_{vision}=-\sum_{\mathrm{k}=1}^{\mathrm{N}}(\frac{1}{T}\sum_{F\in\mathcal{F}}\operatorname{CLIP}\left(F,s_{k}\right))\log\left(\mathrm{q}\left(\mathrm{w}_{k}\right)\right)$ (6)
@@
where $T$ is the total number of sampled video frames. CLIP calculates the average matching score between each sampled frame $F$ in $\mathcal{F}$ and $s_{k}$.
@@
$\mathcal{L}_{\text{language}}=\text{CE}\left(\mathrm{q}(w_{i+1}),\mathrm{q}(\hat{w}_{i+1})\right)$ (7)
$\hat{w}_{i+1}=\operatorname{PLM}(w_{hard},s_{i})$ (8)
@@
3.3 Inference
In our zero-shot setting, we directly perform test-time learning, where only text corpus of training split are used for retrieving caption candidates. The retrieved captions are used to generated soft targets as mentioned in Sec. 3.2. The learnable tokens are optimized by the total loss:
$\mathcal{L}_{total}=\lambda_{l}\mathcal{L}_{language}+\lambda_{s}\mathcal{L}_{retrieval}^{S}+\lambda_{w}\mathcal{L}_{retrieval}^{W}+\lambda_{v}\mathcal{L}_{vision}$ (9)